OpenStudy (anonymous):

what is 114 times 17 over 38 times 1 over 9

OpenStudy (anonymous):

just multiply across 144/1 * 17/38 144*17 = 2448 1*38 = 38 2448/38= 1224/19 now 1224/19 * 1/9 1224*1=1224 19*9 = 171

OpenStudy (anonymous):

1224/171= 136/19 thats it..